If chocolate, nuts and caramel are to be mixed in the ratio 3:5:7 respectively, and 5 tons of chocolate are available, how many tons of the mixture can be made? (Assume there is enough nuts and caramel to use all the chocolate.)
Found 2 solutions by checkley77, edjones:
Answer by checkley77(12844)

3x=5,000
x=5,000/3
x=1,666.67
5*1,666.67=8,333.33 tons of nuts.
7*1,666.67=11,666.67 tons of caramel.
Answer:
(5000):(8,333.33):(11,666.67) tons.

Answer by edjones(8007)

3+5+7=15
5/3 * 15 = 25 tons
.
Ed